OpenStudy (anonymous):

Consider the function. F(x) = (1/5)x - 4 a.) Find the inverse of f(x) and name it g(x). Show and explain your work.

OpenStudy (anonymous):

You basically switch x and y, then solve for y to find an inverse Start: y = .2x - 4 Switch x and y: x = .2y - 4 x+4 = .2y 5x + 20 = y And thats it!

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

You bet!